OpenStudy (anonymous):

Which of the following is not a way to represent the solution of the inequality 2(x − 2) < or equal to 2 A number line with a closed circle on 3 and shading to the right. A number line with a closed circle on 3 and shading to the left. 3 > or equal to x x < or equal to 3

OpenStudy (jiteshmeghwal9):

\[2+4\le2\]\[2x \le 2-4\]\[2x \le 6\]\[x \le \frac{6}{2}\]\[x \le 3\]

OpenStudy (jiteshmeghwal9):

the first step is 2x+4 le 2*